HUBBLE SPACE TELESCOPE DAILY REPORT #2227 PERIOD COVERED: 0000Z (UTC) 10/09/98 - 0000Z (UTC) 10/13/98 Daily Status Report as of 286/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Eight Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD/MA2/WF/PC-2/NIC/1/2/3 8076 (HDF-South Observations) The WF/PC-2, The Space Telescope Imaging Spectrograph (MA2 and CCD) and the Near Infrared Camera (Cameras One, Two and Three) all were used to make deep observations of the Hubble Deep Field - South (HDFS-22333). This program is designed to provide deep observations of a high-galactic latitude field in the southern continuous viewing zone. The STIS field is centered on a bright QSO at z=2.25. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Six Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. During the fourth iteration of this proposal, the acquisition default to fine lock backup described in 2.1 and HSTAR 6709 occurred. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.4 Completed S/C 5582 (FOC/48 Monthly Activation (Camera Section Only)) This proposal was executed as scheduled, and no anomalies were noted. 1.5 Completed Eleven Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Five Sets of STIS/MA2 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark levels in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.7 Completed Six Sets of STIS/CCD 8071 (HDF-South Flanking Fields) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the distant object SAO255267 as part of the HDF-South flanking field study. This program is designed to provide coverage around the HDF-South field and to provide a deep optical images of the NICMOS HDF-South field. The images were completed as planned, and no problems were reported. 1.8 Completed FOC/96 7923 (FOC F/96 Turnon) The Faint Object Camera (f/96) was used to take an internal flat and a dark to ensure that the 30-day limit for the FOC is not exceeded. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed NIC/1 7892 (Narrow-band Imaging of BAL QSOs) The Near Infrared Camera (Camera One) was used to make narrow band observations of the BL-Lac (BAL) QSO Q0021-0213. The detection of extended emission-line gas structures in BAL QSOs would offer another means of investigating the BAL QSO phenomenon as well as establishing a new probe of material in galaxies at high redshift. The observations were completed as planned, and no problems were reported. 1.10 Completed WF/PC-2 7622 (WF/PC-2 Cycle 7 Internal Monitor) The WF/PC-2 was used to make a set of Cycle 7 internal monitoring calibration measurements. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. A bias calibration and an internal flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 7476 (Sakurai's Nova-Like Object: Real-Time Monitoring of a Stellar Thermal Pulse) The WF/PC-2 was used to make images of Sakurai's nova-like object in order to do real-time monitoring of a stellar thermal pulse. The eruption of Sakurai's nova-like variable in Sagittarius presents a "once-in-a-lifetime" opportunity for real-time observations of a star undergoing a final helium thermal pulse. The observations were completed as planned, and no anomalies were reported. 1.12 Completed Three Sets of NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed Ten Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.14 Completed NIC/1/2/3 7886 (NICMOS Snap Shot Survey of Early-Type Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the early-type galaxy NGC-821 as part of a snap-shot broad-band imaging survey. Since almost all early-type galaxies harbor dust, these images which are less sensitive to extinction and will provide superior light profiles for dynamical studies of the cores of these galaxies. The observations were completed as scheduled, and no problems were reported. 1.15 Completed WF/PC-2 6554 (Extragalactic Globular Cluster Systems) The WF/PC-2 was used to make observations of a extragalactic globular cluster system which is located in galaxy NGC-7619. This is part of a survey which will cover a range in galaxy luminosity, morphological type, environment, galactic cluster specific frequency and Galactic cluster mean metallicity. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed Eight Sets of NIC/2/3 8082 (NICMOS pure parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed Two Sets of STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.18 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star V1581-Cyg. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.19 Completed NIC/2 7897 (Probing Planetary Formation Around Main-Sequence Stars: A Snapshot Survey) The Near Infrared Camera (Camera Two) was used to make snapshort survey images of the main sequence star HR7012 probing for planetary formation. The IR excess main-sequence stars host cool circumstellar dust disks, which are most likely replenished by the debris of unseen planetesimals. The observations were completed as scheduled, and no problems were noted. 1.20 Completed WF/PC-2 7628 (WF/PC-2 Cycle 7 Photometric Characterization) The WF/PC-2 was used to make observations of globular cluster NGC-2100 as part of a photometric characterization calibration observation. This is to determine whether any changes in the zeropoint, the spatial dependence of the zeropoint, or contamination rates have occurred. The observation was executed as scheduled, and no anomalies were reported. 1.21 Completed NIC/1/2/STIS/CCD/MA1 7565 (Comparison of Protoplanetary Disks) The Near Infrared Camera (Cameras One and Two) and the Space Telescope Imaging Spectrograph (CCD and MA1) were used to make observations of the protoplanetary system around star HD-163296. This is to search for planets and disk structures around nearby bright stars, comparing an example of a solar type and a hotter star, each with a main sequence and a pre-main sequence star, using broad band coronagraphic imaging using the STIS CCD. Multiple roll angles were used to minimize systematic errors and image artifacts. The observations were executed as scheduled, and no anomalies were noted. 1.22 Completed Five Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.23 Completed NIC/1/2/3 7854 (The Onset of Axisymmetry in Proto-Planetary Nebulae) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the axissysmitry of the proto-planetary nebula CIT6. The high-resolution imaging and multicolor polarimetric capabilities of NICMOS comprise a superb combination for studying the processes by which evolved (approximately one solar mass) stars shed their outer layers to form planetary nebulae. The observations were completed as planned, and no problems were reported. 1.24 Completed NIC/1/2 8050 (APM 08279+5255 - An Ultraluminous BAL Quasar at z=3.87?) The Near Infrared Camera (Cameras One and Two) was used to make observations of the ultraluminous BLAC object APM08279+5255. During a survey for distant carbon stars in the Galactic Halo this object was discovered. It was an extremely bright, R=15.2 broad absorption line quasar at a redshift of about z=3.87. During this proposal, the acquisition default to fine lock backup described in 2.1 and HSTAR 6709 occurred. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.25 Completed NIC/1/2/3 7842 (A High-Resolution Proper Motion Study of the Ionized Gas Near Sgr A*) The Near Infrared Camera (Cameras One, Two and Three) was used to make high resolution proper motion observations of the ionized gas near the black hole candidate GCPAA1. The data used to derive proper motions will be combined with observations from approved proposal 7320 to increase the time baseline of the proper motion measurement. The observations were executed as scheduled, and no anomalies were noted. 1.26 Completed NIC/1/2/3 7268 (Nuclear Brightness Profiles of Merger Remnants; Do Mergers Form Ellipticals?) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the elliptical galaxy NGC-34. Many elliptical galaxies might have formed from mergers of disk galaxies. Numerical simulations yield elliptical-like remnants and observed merger remnants have R^1/4 surface brightness profiles. However, while the global properties of merger remnants and ellipticals are similar, it is unknown whether this is true for their nuclear properties as well. The observations were executed as scheduled, and no anomalies were noted. 1.27 Completed Two Sets of S/C 7954 (FGS-1R Optimization) Desorption and gravity release effects have modified the morphology and modulation of the FGS1R Transfer Function. This test will accomplish a re-optimization of the performance of FGS1R by adjusting the FF3 mirror until the S curve at the center of the field of view is comparable to the optimized S curve obtained during SMOV. This test requires special data downlinks and uplinks for real-time commanding of the FF3. The test concluded with no reported problems. 1.28 Completed Two Sets of NIC/1/2/3 7239 (Surface Compositions of Kuiper Belt Objects) The Near Infrared Camera (Cameras One, Two, and Three) was used to make observations of the Kuiper Belt object Pholus in order to study its surface composition. The surfaces of Kuiper Belt Objects contain a record of formative and dynamical processes in the outer solar system. The observations were executed as scheduled, and no anomalies were noted. 1.29 Completed NIC/1/2/3 7898 (Snapshot Survey of Proto-planetary Nebulae) The Near Infrared Camera (Cameras One, Two and Three) was used to make snapshot survey observations of the protoplanetary nebula OH53.8+20. This is to help discern the structure of the nebulae and the evolutionary sequence that produces axially symmetric planetary nebulae. The observations were executed as scheduled, and no anomalies were noted. 1.30 Completed Six Sets of NIC/1/2/3 7963 (DARKs: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.31 Completed NIC/2 7329 (The Nature of the Damped LyAlpha Absorbers -- A New Study of Young Galaxies) The Near Infrared Camera (Camera Two) was used to make observations of the young galaxy Q0809+4822 in order to study the nature of its damped Lyman-Alpha Absorbers (DLAs). The DLA line systems, occasionally found in quasar spectra, provide one of only a few practical methods of detecting and studying significant numbers of galaxies at high redshifts. The observations were executed as scheduled, and no anomalies were noted. 1.32 Completed NIC/1/2/3 7179 (Search For Volcanic Activity on Encleadus) The Near Infrared Camera (Camera One, Two and Three) was used to search for volcanic activity on Saturn's moon Enceladus. It is widely believed that Enceladus must be the source of material in Saturn's E ring, because the number density of ring particles is highest at precisely the orbit of Enceladus. The observations were completed as planned, and no anomalies were reported. 1.33 Completed STIS/CCD 7672 (STIS CCD Sensitivity Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard star AGK+81D266 in order to monitor the sensitivity of each CCD grating mode to detect any change due to contamination or other causes. The observations were completed as scheduled, and no anomalies were noted. 1.34 Completed STIS/CCD 7404 (STIS Spectroscopy of Markarian 78) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the galaxy Markarian 78 in order to study spectrographically the interaction between its jets and the interstellar medium in the central few kpc of this Seyfert class galaxy. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.35 Completed FOC/96 7686 (FOC Cycle 7 PSF Monitoring - Version 2) The Faint Object Camera (f/96) was used to measure the point spread function (PSF) by making observations of standard star BPM16274. Data will be taken four times during the cycle. The observations were executed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 33 Successful : 33 The acquisition at 284/0440Z defaulted to fine lock backup on FGS-3 only due to a walkdown failure on FGS-1. HSTAR 6709 was written. The proposals detailed in 1.3 and 1.24 may have been affected. Scheduled Reacquisitions : 31 Successful : 31 2.2 FHST Updates: Scheduled : 70 Successful : 70 2.3 Operations Notes: Realtime operations was transferred to the SAMS backup facility at 282/2039Z to allow for scheduled PDU power work. Operations returned to the SEER at 284/0645Z. There were no problems with either transition. Operations functions were performed nominally at SAMS. Using ROP SR-1, the SST EDAC error counter was reset ten times. As documented in HSTARs 6707 and 6708, there were NICMOS single bit errors at 283/090556Z and at 284/004453Z. In each case the NICMOS internal log was dumped, and the ground system limit was changed, the last time to 23. The STIS MCE-2 reset at 284/050306Z during the entry of an SAA. ROP NS-12 was executed to clear the error message. Recovery was performed via SMS commanding at the time of the next MAMA-2 high voltage on commanding. The STIS MCE-1 reset at 284/152047Z. ROP NS-12 was performed to clear the error counter. Rocovery of the MAMA-1 occuurred via SMS commanding at the time of the next high voltage on event (284/154106Z). The FGS-1R fold flat #3 mirror was moved at 285/0111Z. The objective of this move was to return the pupil decenter to the optimum position obtained on day 128 to maximize the astrometric sensitivity and to concurrectly update the onboard alignment transformation coefficients. New FGS table loads were uplinked at 285/0151Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH